Suppose that you have a supply of a 25% solution of alcohol and a 75% solution of alcohol. How many quarts of each should be mixed to produce 70 quarts that is 35% alcohol?
Answer by stanbon(75887)

Suppose that you have a supply of a 25% solution of alcohol and a 75% solution of alcohol. How many quarts of each should be mixed to produce 70 quarts that is 35% alcohol?
---
Equation:
alcohol + alcohol = alcohol
0.25x + 0.75(70-x) = 0.35*70
Multiply thru by 100 to get:
25x + 75*70 - 75x = 35*70
-50x = -40*70
x = 56 quarts (amount of 25% solution needed in the mix)
70-x = 14 quarts (amount of 75% solution needed in the mix)
==========================
Cheers,
Stan H.